exam question...8(5r-3) is less than or = to 4(10r+1).....thanks
Answer by stanbon(75887)

exam question...8(5r-3) is less than or = to 4(10r+1).....thanks
-------
8(5r-3) <= 4(10r+1)
------
40r - 24 <= 40r + 4
------
-24 <= 4
-----
The inequality is true for all values of "r".
Cheers,
Stan H.
=================